OpenStudy (anonymous):

medal! pls help solve 4/x+2 = 2/x

OpenStudy (anonymous):

\[\frac{4}{x+2}=\frac{2}{x}\] \[\frac{x+2}{4}=\frac{x}{2}\]

OpenStudy (anonymous):

for x and determine if the solution is extraneous or not.

OpenStudy (anonymous):

sorry i forgot to post that part..

OpenStudy (anonymous):

flip it,the solve \[x+2=2x\]

OpenStudy (anonymous):

what do you mean by flip it O.o

OpenStudy (anonymous):

*then

OpenStudy (anonymous):

what do you mean by "what do you mean"?

OpenStudy (anonymous):

how do you flip it?

OpenStudy (anonymous):

\[\frac{4}{x+2}=\frac{2}{x}\] take the "reciprocal" of both sides

OpenStudy (anonymous):

you get \[\frac{x+2}{4}=\frac{x}{2}\]

OpenStudy (swamijatin2):

i was also not understanding this step

OpenStudy (anonymous):

since 4 is twice 2 that means \[x+2=2x\]

OpenStudy (anonymous):

im really bad at math lol. these were the answer choices x = -2, extraneous x = -2, non-extraneous x = 2, extraneous x = 2, non-extraneous

OpenStudy (anonymous):

so it would be 2?

OpenStudy (anonymous):

pick one, you have a one in four chance of getting it right yes, it is 2\[x+2=2x\\x=2\]